Circular No. 7269

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999eb AND GRB 991002 R. Terlevich and A. Fabian, Institute of Astronomy, Cambridge; and

M. Turatto, Osservatorio Astronomico, Padova, write: “We note the approximate coincidence between the positions of the bright type-IIn supernova 1999eb in NGC 664 (IAUC 7268) and the γ-ray burst GRB 991002. The position of GRB 991002 was given in the Huntsville Administrative Report (BATSE Trigger No. 7784; cf. http://gcn.gsfc.nasa.gov/gcn/ batse grbs.html) as α = 1^{h}40^{m}.6, δ = +3^{o}45′ (equinox 2000.0), with a 1σ error circle of 2^{o}.2, and its offset with respect to the optical position of SN 1999eb is ∼ 0^{o}.9. Interestingly, the data suggest that the optical flare started around Sept. 19, and by Oct. 2 the supernova was possibly still on the rising branch of the light curve. Type-IIn supernovae are rare events (up to six detected per year), and there is ∼ 1 GRB detected per day. Among the previous reported coincidences between supernovae and GRBs, at least one supernova was of type IIn (SN 1997cy with GRB 970514); in this case, the chance coincidence (a posteriori) was 0.007 percent (Germany et al. 1999,

http://xxx.lanl.gov/abs/astro-ph/?9906096). Because SN 1999eb is

bright and may be close to its maximum, it is crucial to obtain good multifrequency data over the next few months. This type of event is expected to be bright from hard x-rays to radio (Fabian and Terlevich 1996, MNRAS 280, L5; Aretxaga et al. 1999, ibid. 309, 343), so that x-ray spectra and radio monitoring are needed.”

SUPERNOVA 1999ec IN NGC 2207 S. Jha, P. Garnavich, P. Challis and R. Kirshner, Harvard-Smithson-

ian Center for Astrophysics, report that a spectrum of SN 1999ec, obtained by P. Berlind on Oct. 4.5 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), reveals it to be a type-Ib supernova past maximum light. The spectrum exhibits prominent He I absorption and weaker Fe II features. The recession velocity of NGC 2207, measured via narrow emission lines from a superimposed H II region, is 2740 km/s.

SUPERNOVA 1999do IN MARKARIAN 922 CCD magnitude estimates by L. Kiss, K. Sárneczky, and G. Szabó,

Piszkéstető, Konkoly Observatory: Sept. 23.84 UT, 16.7 ± 0.3.

1999 October 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green